# DebiasDiff: Debiasing Text-to-image Diffusion Models with Self-discovering Latent Attribute Directions

### v1

@@ Abstract
While Diffusion Models (DM) exhibit remarkable performance across various image generative tasks, they nonetheless reflect the inherent bias presented in the training set. As DMs are now widely used in real-world applications, these biases could perpetuate a distorted worldview and hinder opportunities for minority groups. Existing methods on debiasing DMs usually requires model re-training with a human-crafted reference dataset or additional classifiers, which suffer from two major limitations: (1) collecting reference datasets causes expensive annotation cost; (2) the debiasing performance is heavily constrained by the quality of the reference dataset or the additional classifier. To address the above limitations, we propose DebiasDiff, a plug-and-play method that learns attribute latent directions in a self-discovering manner, thus eliminating the reliance on such reference dataset. Specifically, DebiasDiff consists of two parts: a set of attribute adapters and a distribution indicator. Each adapter in the set aims to learn an attribute latent direction, and is optimized via noise composition through a self-discovering process. Then, the distribution indicator is multiplied by the set of adapters to guide the generation process towards the prescribed distribution. Our method enables debiasing multiple attributes in DMs simultaneously, while remaining lightweight and easily integrable with other DMs, eliminating the need for re-training. Extensive experiments on debiasing gender, racial, and their intersectional biases show that our method outperforms previous SOTA by a large margin. Our code will be available at https://github.com/leigest519/DebiasDiff.
@@ 1 Introduction
State-of-the-art Text-to-Image Diffusion Models (DMs) such as Stable Diffusion [26], DALL-E 3 [20] and Imagen [27] have demonstrated remarkable performance in generating high-quality images. With the rapid development of DMs, an increasing number of individuals and corporations are choosing to utilize them to serve their own purposes. For instance, Stable Diffusion v1.5 has been downloaded over 8 million times on the Huggingface repository, and Midjourney is used by over a million users [5]. However, existing DMs have been found to generate biased content across various demographic factors, such as gender and race [13], which could have harmful effects on society when these models are implemented in real-world applications.
@@
In Figure 1, we randomly generate several images of two occupations using Stable Diffusion v2.1. Given the prompt of ‘A photo of a CEO’, the generated images predominantly depict male figures, reinforcing the stereotype that leadership roles, such as CEOs and doctors, are male-dominated. On the contrary, when the prompt is ‘A photo of a nurse’, the majority of generated images depict female figures, reflecting the bias that supportive roles are traditionally associated with women. Regarding racial bias, we randomly generate 1000 images using Stable Diffusion v2.1 with the prompt ‘A photo of a worker’. The statistic of 1000 images depicted in Figure 3 shows a strong bias in racial representation, with White individuals making up 71% of the total, while minority groups like Middle Eastern, Latino, Black, and Indian each account for only 3-4%. This bias in DM, produce less accurate or fair results for underrepresented populations. We further investigate such bias situation across across different versions of DMs. We randomly generate 1000 images with the prompt of ‘A photo of a CEO’ using five versions of DMs: Stable Diffusion v1.3 [21], v1.4 [22], v.1.5 [23], v2.0 [24]and v2.1 [25]. Figure 3 demonstrates that gender bias exists across all versions of DMs.
@@
Figure 2: Racial bias in randomly generated 1000 images with the prompt ‘A photo of a worker’ using Stable Diffusion v2.1.
@@
Figure 3: Gender bias in randomly generated 1000 images with the prompt ‘A photo of a CEO’ using Stable Diffusion v1.3, v1.4, v1.5, v2.0, and v2.1.
@@
This bias arises from two main factors. First, the training data sourced from the web is inherently biased. Second, the bias is partly inherited from the CLIP [19] model used in the generation process. Biases in the generated data are even more pronounced in large text-to-image DMs, where models often produce content that associates specific genders with particular professions. Several work has attempted to mitigate the bias in DMs by re-training the model [30] or training-free approaches [18, 8, 17]. In training-based methods, Shen et al. [30] propose to manually curate a reference dataset and match the distribution of generated images with that of the reference dataset. Among training-free approaches, Parihar et al. [18] exploit the rich demographic information embedded in the latent features of the denoising U-Net to guide the generation. Although this method avoids re-training DMs, it depends on training an MLP-based Attribute Distribution Predictor using pseudo labels generated from existing attribute classifiers. This reliance on accurate attribute classifiers for training $h$-space classifiers significantly limits its debiasing performance. Additionally, Gandikota et al. [8] and Orgad et al. [17] employ closed-form editing to adjust concepts within DMs without re-training. However, training-based methods heavily depend on gathering annotated reference datasets, which are both expensive and constrained by the dataset’s quality. Despite the implementation efficiency of training-free methods, they tend to be less effective than training-based approaches.
@@
To address these limitations, we propose DebiasDiff, a plug-and-play method that achieves debiased image generation (Figure 1) by automatically learning attribute latent directions, mitigating the reliance on reference datasets. DebiasDiff is composed of two components: a set of attribute adapters and a distribution indicator. Each adapter is trained to learn an attribute-specific latent direction, optimized through noise composition in a self-discovering manner, i.e., our method automatically learns attribute latent directions without relying on a labeled reference dataset. Through noise composition, our method explores and optimizes attribute directions directly from the model’s latent space, uncovering patterns without external supervision. At inference stage, the distribution indicator is then applied to select attribute-specific adapters, guiding the generative process towards the desired distribution. We comprehensively evaluate the effectiveness of our approach in debiasing gender, racial, and intersectional biases using occupational prompts. Experimental results demonstrate that our method not only achieves state-of-the-art performance in single and multiple attribute debiasing tasks but also preserves the generation quality of DM. Furthermore, we show that once DebiasDiff is trained on one diffusion model, it can be seamlessly integrated into other models without re-tuning. Thanks to its strong transferability and plug-and-play functionality, our method offers a practical solution for both individual users and organizations, facilitating the responsible use of diffusion models in future applications.
@@
  • We propose DebiasDiff, a novel method for debiasing DMs by learning attribute latent directions in a self-discovering manner, eliminating the reliance on the reference dataset or classifier, and thus significantly reduce the cost.
  • DebiasDiff is lightweight, plug-and-play and shows good transferability across different DMs, making it more convenient to deploy in the real world.
@@ 2 Related Work
Debiasing Diffusion Models by retraining. Before DMs, previous approaches mainly focus on debiasing GANs by assuming access to the labels of sensitive attributes and ensure no correlation exists between the decision and sensitive attributes [16, 35, 33, 28, 36, 3, 31, 32]. More recently, regarding DMs, Shen et al. [30] propose a distributional alignment loss to guide the characteristics of the generated images towards target distribution and use adjusted finetuning to directly optimize losses on the generated images. This method requires a reference training dataset to complete the retraining process, whereas our method does not need such reference dataset, which largely reduces annotation costs. Li et al. [11] find a interpretable latent vector for a given concept, and achieve debiasing by manipulating the h-space inside DMs during inference stage.
Debiasing Diffusion Models without training. Parihar et al. [18] propose Distribution Guidance (DG), which guides the generated images to follow the prescribed attribute distribution. Although DG does not require retraining DMs, it requires training an Attribute Distribution Predictor (ADP) that maps the latent features to the distribution of attributes. Since ADP is trained with pseudo labels generated from existing attribute classifiers, the performance of DG is largely constrained by the accuracy of attribute classifiers. Gandikota et al. [8] and Orgad et al. [17] use closed-form editing approach to edits concepts inside DM without training. Despite the efficiency, these methods are less effective than training-based approachs.
@@ 3 Preliminary
Latent Diffusion Models (LDMs)
[26], also known as Stable Diffusion (SD), perform the diffusion process within the latent space. During training, noise is added to the encoded latent representation of the input image $x$, resulting in a noisy latent code $z_{t}$ at each time step $t$.
In the pretraining stage, an autoencoder framework is employed to map images into a lower-dimensional latent space via an encoder: $z=\mathcal{E}(x)$. The decoder then reconstructs images from these latent codes: $x\approx\mathcal{D}(\mathcal{E}(x))$. This process ensures that the latent space retains the essential semantic information of the image.
The training objective of the diffusion model in the latent space is given by:
@@
where $\epsilon$ is Gaussian noise sampled from a normal distribution $\mathcal{N}(0,1)$, $\epsilon_{\theta}$ is the denoising network, and $c$ represents any conditioning embeddings (e.g., text or class labels).
At the inference stage, a latent code $z_{T}$ is sampled from Gaussian noise at the initial timestep $T$. The denoising network $\epsilon_{\theta}$ is then applied iteratively to remove the noise over several steps, generating a denoised latent representation $z_{0}$. Finally, the pretrained decoder reconstructs the image from the denoised latent code: $\hat{x}_{0}\approx\mathcal{D}(z_{0})$, where $\hat{x}_{0}$ is the generated output image.
Classifier-free Guidance
[9] aim to modulate image generation by steering the probability distribution towards data that is more probable according to an implicit classifier $p(c\mid z_{t})$. It operates at inference phase and the model is jointly trained on both conditional and unconditional denoising tasks. During inference, both the conditional and unconditional denoising scores are derived from the model. The final score $\tilde{\epsilon}_{\theta}(z_{t},c,t)$ is then adjusted by weighting the conditioned score more heavily relative to the unconditioned score using a guidance scale $\alpha>1$.
@@ 4 Method
Figure 4: Overview of DebiasDiff’s training pipeline. Attribute-specific adapters ($M$) are attached to the cross-attention layers in the denoising UNet. Target group and attribution direction are fed into the DM for composing the noise predictions (Eq. 5), which is used as self-discovering attribute direction guidance to optimize the adapters.
Inspired by recent work that fine-tune diffusion models for erasing concepts [6] and controlling the level of concepts [7] (e.g., ages, size of eyes etc), we aim at reducing the bias in the DM by attaching and learning a set of light-weight adapters, each of which represents a category of an attribute (e.g., female of gender), guiding the DM towards an attribute latent direction. Unlike previous work that relies on additional reference datasets and has to finetune the whole DM [30], we instead optimize the attached adapters via noise composition through a self-discovering process (detailed in Section 4.2). This significantly reduces the cost in computation and data. During the inference stage, given a predefined target distribution (e.g., uniform), we introduce a distribution indicator implemented by a gating function to select one corresponding adapter which will be attached to the DM for generating image. In this way, the set of generated images will follow the predefined target distribution, and thus are not biased to some categories of an attribute (if the predefined target distribution is uniform). Our training and inference diagrams are illustrated in Figure 4 and Figure 5. In the following sections, we start elaborating our method in single attribute settings and then extend it to more general ones.
Figure 5: Overview of DebiasDiff’s inference pipeline. The distribution indicator is generated according the prescribed distribution. Then, it is multiplied by the set of attribute adapters matrices to select attribute matrix adapter. The selected adapter is integrated to the DM with no overhead, guiding the generation towards the prescribed distribution.
@@ 4.1 Attaching Light-weight attribute-specific Adapters
Specifically, for the $i$-th cross-attention layer parameterized by $\bm{W}_{i}\in\mathbb{R}^{m\times n}$ in the denoising U-Net, we attach an adapter to the layer to guide the attribute towards a certain category. The adapter consists of two 1-dim vectors: $\bm{p}\in\mathbb{R}^{m}$ and $\bm{q}\in\mathbb{R}^{n}$. The forward process of the $i$-th cross-attention layer would be updated from $\bm{y}_{i}=\bm{W}_{i}\bm{x}_{i}$ to $\bm{y}_{i}=\bm{W}_{i}\bm{x}_{i}+(\bm{q}_{i}^{T}\bm{x}_{i})\cdot\bm{p}_{i}$. $\bm{x}_{i}\in\mathbb{R}^{n}$ and $\bm{y}_{i}\in\mathbb{R}^{m}$ represent the input and output of the layer, and superscript $T$ indicates transposition. Thus, adapters in all ($r$) cross-attention layers for an attribute category $d$ are:
$M_{d}=\bm{Q}^{T}\bm{P},$ (3)
where $\bm{Q}=\left[\bm{q}_{1},\bm{q}_{2},\ldots,\bm{q}_{r}\right],\bm{P}=\left[\bm{p}_{1},\bm{p}_{2},\ldots,\bm{p}_{r}\right].$ Each column vector in $\bm{Q}$ and $\bm{P}$, we pad 0 to their end if their dimensions are not the same.
@@ 4.2 Optimizing Adapters via self-discovering process
Given a target group $g_{t}$ (for example, ‘CEO’) and model $\theta$, we want to optimize the adapters such that the model attached with the adapters generate image $X$ towards certain attribute category $d$ (for example, ‘male’ or ’female’ ) when conditioned on $g_{t}$:
@@
where $P_{\theta}(X|g_{t})$ represents the distribution generated by the original model when conditioned on $g_{t}$, and $\theta^{*}$ represents the new model equipped with the adapters of $d$. Note that $g_{t}$ can be set to an empty string ‘’ so that the model will be debiased for all possible groups.
@@
Based on Tweedie’s formula [4] and the reparametrization trick of Classifier-free guidance [9], we introduce a time-varying noising process and represent each score (gradient of log probability) as a denoising prediction $\epsilon(X_{t},c_{t},t)$, which leads to our learning objective for adapters of category $d$ of an attribute.
$\begin{split}\epsilon_{\theta^{*}}(X_{t},g_{t},t)\leftarrow\;\epsilon_{\theta}(X_{t},g_{t},t)\;+\eta\left(\epsilon_{\theta}(X_{t},d,t)-\epsilon_{\theta}(X_{t},g_{t},t)\right)\end{split}$ (6)
@@
$\begin{split}\mathcal{L}_{\text{Guidance }}=\mathbb{L}_{x_{t},t}\left[\left\|\epsilon_{\theta^{*}}(X_{t},g_{t},t)-\epsilon_{L}\right\|^{2}\right],\\ \epsilon_{L}=\epsilon_{\theta}(X_{t},g_{t},t)\;+\eta\left(\epsilon_{\theta}(X_{t},d,t)-\epsilon_{\theta}(X_{t},g_{t},t)\right)\end{split}$ (7)
@@
After the optimization, we obtain a set of adapters =$\mathcal{M}=\{\bm{M}_{1},\bm{M}_{2},\ldots,\bm{M}_{t}\}$ for $t$ categories of a given attribute. For instance, $t=2$ for gender bias and $t=4$ for racial bias in our case. As shown in Figure 5, at the inference stage, we introduce a distribution indicator $\bm{h}$. Given a prescribed distribution ${f^{a}_{\theta}}$ (e.g., uniform distribution), we define its Probabilistic Mass Function (PMF) as follows:
@@
Then, we randomly sample an index $k$ from the prescribed distribution ${f^{a}_{\theta}}$. The distribution indicator $\bm{h}\in\mathbb{R}^{t}$ is formulated to reflect the chosen index $k$ as follows:
$\bm{h}_{i}:=\begin{cases}1&\text{if }i=k,\\ 0&\text{if }i\neq k,\end{cases}$ (9)
where $i\in\{1,2,\dots,t\}$, and $k$ is the sampled index based on the distribution ${f^{a}_{\theta}}$. The indicator $\bm{h}$ is a one-hot vector, where the $k$-th element is 1, indicating the sampled outcome, and all other elements are 0. After obtaining the distribution indicator, it is multiplied by the set of trained attribute matrix adapters. Subsequently, the final weight change $\Delta\bm{W}$ is given by:
$\Delta\bm{W}=\bm{h}\cdot\mathcal{M}.$ (10)
And the model is updated as $\bm{W}\leftarrow\bm{W}+\alpha\Delta\bm{W}$, where $\alpha$ is a scaling factor controlling the strength of the guidance.

### v3

# FairGen: Enhancing Fairness in Text-to-Image Diffusion Models via Self-Discovering Latent Directions

@@ Abstract
While Diffusion Models (DM) exhibit remarkable performance across various image generative tasks, they nonetheless reflect the inherent bias presented in the training set. As DMs are now widely used in real-world applications, these biases could perpetuate a distorted worldview and hinder opportunities for minority groups. Existing methods on debiasing DMs usually requires model retraining with a human-crafted reference dataset or additional classifiers, which suffer from two major limitations: (1) collecting reference datasets causes expensive annotation cost; (2) the debiasing performance is heavily constrained by the quality of the reference dataset or the additional classifier. To address the above limitations, we propose FairGen, a plug-and-play method that learns attribute latent directions in a self-discovering manner, thus eliminating the reliance on such reference dataset. Specifically, FairGen consists of two parts: a set of attribute adapters and a distribution indicator. Each adapter in the set aims to learn an attribute latent direction, and is optimized via noise composition through a self-discovering process. Then, the distribution indicator is multiplied by the set of adapters to guide the generation process towards the prescribed distribution. Our method enables debiasing multiple attributes in DMs simultaneously, while remaining lightweight and easily integrable with other DMs, eliminating the need for retraining. Extensive experiments on debiasing gender, racial, and their intersectional biases show that our method outperforms previous SOTA by a large margin. Our code is made publicly avaibable at https://github.com/leigest519/FairGen.
@@ 1 Introduction
State-of-the-art Text-to-Image Diffusion Models (DMs) such as Stable Diffusion [26], DALL-E 3 [20] and Imagen [27] have demonstrated remarkable performance in generating high-quality images. With the rapid development of DMs, an increasing number of individuals and corporations are choosing to utilize them to serve their own purposes. For instance, Stable Diffusion v1.5 has been downloaded over 8 million times on Huggingface, and Midjourney is used by over a million users [5]. However, existing DMs have been found to generate biased content across various demographic factors, such as gender and race [13], which could have harmful effects on society when these models are implemented in real-world applications.
@@
In Figure 1, we randomly generate several images of two occupations using Stable Diffusion v2.1. Given the prompt of “A photo of a CEO”, the generated images predominantly depict male figures, reinforcing the stereotype that leadership roles, such as CEOs and doctors, are male-dominated. On the contrary, when the prompt is “A photo of a nurse”, the majority of generated images depict female figures, reflecting the bias that supportive roles are traditionally associated with women. Regarding racial bias, we randomly generate 1000 images using Stable Diffusion v2.1 with the prompt “A photo of a worker”. The statistic of 1000 images depicted in Figure 3 shows a strong bias in racial representation, with White individuals making up 71% of the total, while minority groups like Middle Eastern, Latino, Black, and Indian each account for only 3-4%. This bias in DM, produce less accurate or fair results for underrepresented populations. We further investigate such bias situation across across different versions of DMs. We randomly generate 1000 images with the prompt of “A photo of a CEO” using five versions of DMs: Stable Diffusion v1.3 [21], v1.4 [22], v.1.5 [23], v2.0 [24] and v2.1 [25]. Figure 3 demonstrates that gender bias exists across all versions of DMs.
@@
Figure 2: Racial bias in randomly generated 1000 images with the prompt “A photo of a worker” using Stable Diffusion v2.1.
@@
Figure 3: Gender bias in randomly generated 1000 images with the prompt “A photo of a CEO” using Stable Diffusion v1.3, v1.4, v1.5, v2.0, and v2.1.
@@
This bias arises from two primary sources. First, the training data, predominantly sourced from the web, contains inherent biases. Second, bias is partially inherited from the CLIP [19] model used in the generation process. In large text-to-image DMs, these biases are even more pronounced, often manifesting in gendered associations with specific professions. Several studies have attempted to mitigate bias in DMs through retraining [30, 36] or approaches that do not require retraining the DM [18, 8, 17]. In training-based methods, Shen et al. [30] propose curating a reference dataset and aligning the distribution of generated images with this dataset. Parihar et al. [18] avoid retraining the DM by leveraging demographic information embedded in the latent features of the denoising U-Net to guide generation. However, their method requires training an additional MLP-based Attribute Distribution Predictor using pseudo labels from existing attribute classifiers. This reliance on accurate attribute classifiers limits its debiasing performance. In contrast, Gandikota et al. [8] and Orgad et al. [17] employ closed-form editing to adjust concepts within DMs without retraining. While training-based methods depend on costly and quality-constrained annotated reference datasets, methods that do not require retraining the DM offer greater implementation efficiency but often exhibit lower debiasing effectiveness.
@@
To address these limitations, we propose FairGen, a plug-and-play method that achieves debiased image generation (Figure 1) by automatically learning attribute latent directions, mitigating the reliance on reference datasets. FairGen is composed of two components: a set of attribute adapters and a distribution indicator. Each adapter is trained to learn an attribute-specific latent direction, optimized through noise composition in a self-discovering manner, i.e., our method automatically learns attribute latent directions without relying on a labeled reference dataset. Through noise composition, our method explores and optimizes attribute directions directly from the model’s latent space, uncovering patterns without external supervision. At inference stage, the distribution indicator is then applied to select attribute-specific adapters, guiding the generative process towards the desired distribution. We comprehensively evaluate the effectiveness of our approach in debiasing gender, racial, and intersectional biases using occupational prompts. Experimental results demonstrate that our method not only achieves state-of-the-art performance in single and multiple attribute debiasing tasks but also preserves the generation quality of DM. Furthermore, we show that once FairGen is trained on one diffusion model, it can be seamlessly integrated into other models without re-tuning. Thanks to its strong transferability and plug-and-play functionality, our method offers a practical solution for both individual users and organizations, facilitating the responsible use of diffusion models in future applications.
@@
  • We propose FairGen, a novel method for debiasing DMs by learning attribute latent directions in a self-discovering manner, eliminating the reliance on the reference dataset or classifier, and thus significantly reduce the cost.
  • FairGen is lightweight, plug-and-play and shows good transferability across different DMs, making it more convenient to deploy in the real world.
@@ 2 Related Work
Debiasing Diffusion Models by retraining. Before DMs, previous approaches mainly focus on debiasing GANs by assuming access to the labels of sensitive attributes and ensure no correlation exists between the decision and sensitive attributes [16, 35, 33, 28, 36, 3, 31, 32]. More recently, regarding DMs, Shen et al. [30] propose a distributional alignment loss to guide the characteristics of the generated images towards target distribution and use adjusted finetuning to directly optimize losses on the generated images. This method requires a reference training dataset to complete the retraining process, whereas our method does not need such reference dataset, which largely reduces annotation costs. Yu et al. [36] enhance fairness in text-to-image synthesis by incorporating reference images. and leverage visual exemplars to more effectively represent attributes that are difficult to describe in words, such as nuanced variations in skin tones. Li et al. [11] find a interpretable latent vector for a given concept, and achieve debiasing by manipulating the h-space inside DMs during inference stage.
Debiasing Diffusion Models without retraining. Parihar et al. [18] propose Distribution Guidance (DG), which guides the generated images to follow the prescribed attribute distribution. Although DG does not require retraining DMs, it requires training an Attribute Distribution Predictor (ADP) that maps the latent features to the distribution of attributes. Since ADP is trained with pseudo labels generated from existing attribute classifiers, the performance of DG is largely constrained by the accuracy of attribute classifiers. Gandikota et al. [8] and Orgad et al. [17] use closed-form editing approach to edits concepts inside DM without training. Despite the efficiency, these methods are less effective than training-based approachs.
@@ 3 Preliminary
Latent Diffusion Models (LDMs) perform the diffusion process within the latent space [26]. During training, noise is added to the encoded latent representation of the input image $x$, resulting in a noisy latent code $z_{t}$ at each time step $t$. In the pretraining stage, an autoencoder framework is employed to map images into a lower-dimensional latent space via an encoder: $z=\mathcal{E}(x)$. The decoder then reconstructs images from these latent codes: $x\approx\mathcal{D}(\mathcal{E}(x))$. This process ensures that the latent space retains the essential semantic information of the image. The training objective of the diffusion model in the latent space is given by:
@@
where $\epsilon$ is Gaussian noise sampled from a normal distribution $\mathcal{N}(0,1)$, $\epsilon_{\theta}$ is the denoising network, and $c$ represents any conditioning embeddings (e.g., text or class labels). At the inference stage, a latent code $z_{T}$ is sampled from Gaussian noise at the initial timestep $T$. The denoising network $\epsilon_{\theta}$ is then applied iteratively to remove the noise over several steps, generating a denoised latent representation $z_{0}$. Finally, the pretrained decoder reconstructs the image from the denoised latent code: $\hat{x}_{0}\approx\mathcal{D}(z_{0})$, where $\hat{x}_{0}$ is the generated output image.
Classifier-free Guidance aims to modulate image generation by steering the probability distribution towards data that is more probable according to an implicit classifier $p(c\mid z_{t})$ [9]. It operates at the inference phase and the model is jointly trained on both conditional and unconditional denoising tasks. During inference, both the conditional and unconditional denoising scores are derived from the model. The final score $\tilde{\epsilon}_{\theta}(z_{t},c,t)$ is then adjusted by weighting the conditioned score more heavily relative to the unconditioned score using a guidance scale $\alpha>1$.
@@ 4 Method
Figure 4: Overview of FairGen’s training pipeline. Attribute-specific adapters ($M$) are attached to the cross-attention layers in the denoising UNet. Target group and attribution direction are fed into the DM for composing the noise predictions (Eq. 5), which is used as self-discovering attribute direction guidance to optimize the adapters.
Inspired by recent work that fine-tune diffusion models for erasing concepts [6] and controlling the level of concepts [7] (e.g., ages, size of eyes), we aim at reducing the bias in the DM by attaching and learning a set of light-weight adapters, each of which represents a category of an attribute (e.g., female of gender), guiding the DM towards an attribute latent direction. Unlike previous work that relies on additional reference datasets and has to finetune the whole DM [30], we instead optimize the attached adapters via noise composition through a self-discovering process (detailed in Section 4.2). This significantly reduces the cost in computation and data. During the inference stage (Section 4.3), given a predefined target distribution (e.g., uniform), we introduce a distribution indicator implemented by a gating function to select one corresponding adapter which will be attached to the DM for generating image. In this way, the set of generated images will follow the predefined target distribution, and thus are not biased to some categories of an attribute (if the predefined target distribution is uniform). Our training and inference diagrams are illustrated in Figure 4 and Figure 5. In the following sections, we start elaborating our method in single attribute settings and then extend it to more general ones.
Figure 5: Overview of FairGen’s inference pipeline. The distribution indicator is generated according the prescribed distribution. Then, it is multiplied by the set of attribute adapters matrices to select attribute matrix adapter. The selected adapter is integrated to the DM with no overhead, guiding the generation towards the prescribed distribution.
@@ 4.1 Attaching Light-weight attribute-specific Adapters
Specifically, for the $i$-th cross-attention layer parameterized by $\boldsymbol{W}_{i}\in\mathbb{R}^{m\times n}$ in the denoising U-Net, we attach an adapter to the layer to guide the attribute towards a certain category. The adapter consists of two 1-dim vectors: $\boldsymbol{p}\in\mathbb{R}^{m}$ and $\boldsymbol{q}\in\mathbb{R}^{n}$. The forward process of the $i$-th cross-attention layer would be updated from $\boldsymbol{y}_{i}=\boldsymbol{W}_{i}\boldsymbol{x}_{i}$ to $\boldsymbol{y}_{i}=\boldsymbol{W}_{i}\boldsymbol{x}_{i}+(\boldsymbol{q}_{i}^{T}\boldsymbol{x}_{i})\cdot\boldsymbol{p}_{i}$. $\boldsymbol{x}_{i}\in\mathbb{R}^{n}$ and $\boldsymbol{y}_{i}\in\mathbb{R}^{m}$ represent the input and output of the layer, and superscript $T$ indicates transposition. Thus, adapters in all ($r$) cross-attention layers for an attribute category $d$ are:
$M_{d}=\boldsymbol{Q}^{T}\boldsymbol{P},$ (3)
where $\boldsymbol{Q}=\left[\boldsymbol{q}_{1},\boldsymbol{q}_{2},\ldots,\boldsymbol{q}_{r}\right],\boldsymbol{P}=\left[\boldsymbol{p}_{1},\boldsymbol{p}_{2},\ldots,\boldsymbol{p}_{r}\right].$ Each column vector in $\boldsymbol{Q}$ and $\boldsymbol{P}$, we pad 0 to their end if their dimensions are not the same.
@@ 4.2 Optimizing Adapters via self-discovering process
Given a target group $g_{t}$ (for example, “CEO”) and model $\theta$, we want to optimize the adapters such that the model attached with the adapters generate image $X$ towards certain attribute category $d$ (for example, “male” or “female” ) when conditioned on $g_{t}$:
@@
where $P_{\theta}(X|g_{t})$ represents the distribution generated by the original model when conditioned on $g_{t}$, and $\theta^{*}$ represents the new model equipped with the adapters of $d$. Note that $g_{t}$ can be set to an empty string “” so that the model will be debiased for all possible groups.
@@
Based on Tweedie’s formula [4] and the reparametrization trick of Classifier-free guidance [9], we introduce a time-varying noising process and represent each score (gradient of log probability) as a denoising prediction $\epsilon(X_{t},c_{t},t)$, which leads to our learning objective for adapters of category $d$ of an attribute:
$\begin{split}\epsilon_{\theta^{*}}(X_{t},g_{t},t)\leftarrow\epsilon_{\theta}(X_{t},g_{t},t)+\eta\left(\epsilon_{\theta}(X_{t},d,t)-\epsilon_{\theta}(X_{t},g_{t},t)\right).\end{split}$ (6)
@@
$\begin{split}\mathcal{L}_{\text{Guidance }}=\mathbb{L}_{x_{t},t}\left[\left\|\epsilon_{\theta^{*}}(X_{t},g_{t},t)-\epsilon_{L}\right\|^{2}\right],\\ \epsilon_{L}=\epsilon_{\theta}(X_{t},g_{t},t)\;+\eta\left(\epsilon_{\theta}(X_{t},d,t)-\epsilon_{\theta}(X_{t},g_{t},t)\right),\end{split}$ (7)
@@
After the optimization, we obtain a set of adapters $\mathcal{M}=\{\boldsymbol{M}_{1},\boldsymbol{M}_{2},\ldots,\boldsymbol{M}_{t}\}$ for $t$ categories of a given attribute. For instance, $t=2$ for gender bias and $t=4$ for racial bias in our case. As shown in Figure 5, at the inference stage, we introduce a distribution indicator $\boldsymbol{h}$. Given a prescribed distribution $f^{a}_{\theta}$} (e.g., uniform distribution), we define its Probabilistic Mass Function (PMF) as follows:
@@
Then, we randomly sample an index $k$ from the prescribed distribution $f^{a}_{\theta}$}. The distribution indicator $\boldsymbol{h}\in\mathbb{R}^{t}$ is formulated to reflect the chosen index $k$ as follows:
$\boldsymbol{h}_{i}:=\begin{cases}1&\text{if }i=k,\\ 0&\text{if }i\neq k,\end{cases}$ (9)
where $i\in\{1,2,\dots,t\}$, and $k$ is the sampled index based on the distribution $f^{a}_{\theta}$}. The indicator $\boldsymbol{h}$ is a one-hot vector, where the $k$-th element is 1, indicating the sampled outcome, and all other elements are 0. After obtaining the distribution indicator, it is multiplied by the set of trained attribute matrix adapters. Subsequently, the final weight change $\Delta\boldsymbol{W}$ is given by:
$\Delta\boldsymbol{W}=\boldsymbol{h}\cdot\mathcal{M},$ (10)
and the model is updated as $\boldsymbol{W}\leftarrow\boldsymbol{W}+\alpha\Delta\boldsymbol{W}$, where $\alpha$ is a scaling factor controlling the strength of the guidance.